# Integrated Push-and-Pull Update Model for Goal-Oriented Effective Communication

### v1

@@ VI-B Results and Discussion
Afterward, we compare the performance provided by model-based agent decisions discussed in Section IV with that of model-free decisions. The latter is based on reinforcement learning (RL), where each agent separately learns to make decisions through direct interaction with the environment. This learning process is modeled under the state (here, observation) spaces, action sets, and rewards according to Section IV-B, without relying on the construction of a predefined model. To derive model-free decisions, we employ a deep Q-network (DQN) and parameterize an approximate value function for every agent within the E-horizon through a multilayer perceptron (MLP), assisted with the experience replay mechanism. The neural network consists of two hidden layers, each with $64$ neurons, and is trained using the adaptive moment estimation (ADAM) optimizer. The default values for the RL setting are taken from [29], except for the learning rate that is $10^{-4}$, and the discount factor assumed $0.75$, as aligned with Table I. In this regard, Fig. 13 depicts the evolution of the average cumulative effectiveness over time for the push-and-pull model and different model-based agent decisions based on the Monte Carlo (MC) estimation method given in Section V and model-free RL-based decisions. As shown, the MC approach exceeds the RL one under all decision policies. However, the performance gap is not significant in the scenario where both agents apply effect-aware policies.
[…]
[scale=0.4]Media/Effect_time_withRL.eps
@@
Figure 13: The evolution of the average cumulative effectiveness for different policies under the MC and RL approaches in the push-and-pull model.
[…]
[scale=0.4]Media/Effect_batchlength_withRL.eps
@@
Figure 14: The E-horizon length’s impact on the effectiveness for different MC- and RL-based policies in the push-and-pull model.
The reason for the better effectiveness of the MC approach could be the small state (observation) space, especially for the modeling at the SA. Additionally, the length of the E-horizon as the training interval may be another factor. In order to analyze the impact of the E-horizon length on the average effectiveness of the effect-aware and effect-agnostic decisions based on the MC and RL approaches, we plot Fig. 14 for the push-and-pull model. We infer that increasing the length of the E-horizon improves effectiveness, but the MC-based policies reach their highest performance after a certain optimal length. This optimal E-horizon length varies depending on the decision policy. For example, in the scenario where both agents utilize the MC approach and apply effect-aware policies, an E-horizon length of around $316$ time slots is needed to ensure accurate enough estimations and achieve the highest average effectiveness. On the other hand, RL-based policies consistently improve across the plotted region. However, the scenarios in which the AA applies effect-aware policies, exhibit a notable challenge—achieving superior performance to MC-based policies demands an extensive E-horizon length. It is also worth mentioning that the reason for the monotonic decrease in the performance of the scenario with Markovian queries based on the MC approach is that the provided effectiveness is limited, and it grows at a slower pace w.r.t. the time interval we sum up the timely effectiveness.
@@ VI-C Lookup Maps for Agent Decisions
[…]
[scale=0.3]Media/Threshold_SA_0.6.eps
[…]
[scale=0.3]Media/Threshold_SA_0.8.eps
Figure 15: The lookup maps for decisions the SA makes based on (a) $C_{\alpha,\rm max}=0.06$ and (b) $C_{\alpha,\rm max}=0.08$.
[…]
[scale=0.26]Media/Threshold_MA_0.6.eps
[…]
[scale=0.26]Media/Threshold_MA_0.8.eps
Figure 16: The lookup maps the AA utilizes for (a) $C_{\beta,\rm max}=0.06$ and (b) $C_{\beta,\rm max}=0.08$.
As the modeled CMDPs in Section IV-B have finite states, we can depict optimal mode-based decisions derived in Algorithm 1 via a multi-dimensional lookup map for each decision-making agent. Figs. 15 and 16 illustrate the maps for the SA and the AA, respectively, under different maximum discounted costs. The number of dimensions in a map relies on the number of elements constructing every state of the relevant CMDP, with each dimension assigned to one element. With the lookup map in hand, an agent can make optimal decisions in each slot based on its current state. When comparing the same maps for different maximum discounted costs, the observation emerges that the more stringent the cost constraint is, the narrower the agent decision boundaries become. Thus, the maps could vary by changes in the parameters’ values or the goals with different target effectiveness grades. It is noteworthy that in Fig. 16, the maps with $\Theta_{n}=1$ represent the pull-based model, while the push-and-pull model converges to the push-based model with $\Theta_{n}\geq 3$ and $2$ in Figs. 16(a) and 16(b), respectively.
We compute an optimal threshold for each element of the state as a decision criterion, given values of the other elements. Let us consider $\Omega_{\alpha}$ and $\Omega_{\beta}$ as the decision criteria at the SA and AA, respectively. To derive the optimal decision $\alpha_{n}^{*}$ in the $n$-th slot, there are two alternative ways to define the criterion:
@@
    $\alpha_{n}^{*}=\mathbbm{1}\big{\{}i\geq\Omega_{\alpha}(\hat{E}_{n})~{}|~{}\hat{E}_{n}\big{\}}.$ (33)
@@
    $\alpha_{n}^{*}=\mathbbm{1}\big{\{}\hat{E}_{n}\geq\Omega_{\alpha}(v_{n})~{}|~{}v_{n}\big{\}}.$ (34)
@@ VII Conclusion
We investigated decision-making for enhancing the effectiveness of updates communicated in the end-to-end status update system based on a push-and-pull communication model. To this end, we considered that the SA observes an information source and generates updates, and its transmission controller decides whether to send them to the AA or not. On the other side, the AA is responsible for acting based on the updates that are successfully received and the raised queries to accomplish a subscribed goal at the endpoint. After defining a GoE metric, we formulated the decision problem of finding optimal effect-aware policies that maximize the expected discounted sum of the update’s effectiveness subject to cost constraints. Using the dual problem, we cast it to a CMDP solved separately for each agent based on different model components and proposed an iterative algorithm to obtain the decision policies. Our results established that the push-and-pull model, on average, outperforms the push- and pull-based models in terms of energy efficiency and effectiveness, respectively. Furthermore, effect-aware policies at both agents significantly enhance the effectiveness of updates with a considerable difference in comparison to those of periodic and probabilistic effect-agnostic policies used at either or both agent(s). Finally, we proposed a threshold-based decision policy complemented by a tailored lookup map for each agent that employs effect-aware policies. Future works could explore multi-SA scenarios with varying sensing abilities and realizations’ time-dependent importance.

### v2

@@ VI-B Results and Discussion
Afterward, we compare the performance provided by model-based agent decisions discussed in Section IV with that of model-free decisions. The latter is based on reinforcement learning (RL), where each agent separately learns to make decisions through direct interaction with the environment. This learning process is modeled under the state (here, observation) spaces, action sets, and rewards according to Section IV-B, without relying on the construction of a predefined model. To derive model-free decisions, we employ a deep Q-network (DQN) and parameterize an approximate value function for every agent within the E-horizon through a multilayer perceptron (MLP), assisted with the experience replay mechanism. The neural network consists of two hidden layers, each with $64$ neurons, and is trained using the adaptive moment estimation (Adam) optimizer. The default values for the RL setting are taken from [29], except for the learning rate that is $10^{-4}$, and the discount factor assumed $0.75$, as aligned with Table I. In this regard, Fig. 13 depicts the evolution of the average cumulative effectiveness over time for the push-and-pull model. The graph compares the performance achieved by model-based agent decisions with that of model-free (RL-based) ones. As shown, the model-based approach outperforms the model-free approach across all decision policies. However, the performance gap is less significant when both agents apply effect-aware policies. Considering the CMDP modeled at both agents, several factors may contribute to the higher effectiveness of the model-based approach, such as the relatively small state space, especially for the modeling at the SA.
@@
Figure 13: Temporal evolution of average cumulative effectiveness under model-based and model-free (RL-based) policies in the push-and-pull model.
Additionally, the length of the E-horizon potentially influences the performance and complexity of both decision-making approaches. To analyze this impact, Fig. 14 demonstrates how the E-horizon length affects the average effectiveness of effect-agnostic and effect-aware policies under both model-based and model-free approaches for the push-and-pull model. The results show that increasing the E-horizon length enhances performance, reaching its peak at an optimal E-horizon length. This optimal length value varies depending on the policy and approach employed. For instance, when both agents employ effect-aware policies, achieving the highest effectiveness requires approximately $10^{2.9}$ and $10^{3.7}$ slots for the model-based and model-free approaches, respectively, with convergence accuracy specified in Table I. In this context, the required E-horizon length serves as a measure of the complexity of the corresponding algorithm for each approach. For the above example, the model-based algorithm (Algorithm 1) achieves a complexity that is $6.31$ times lower compared to the model-free DQN algorithm.
@@
Figure 14: Impact of E-horizon length on the effectiveness of model-based and model-free (RL-based) policies in the push-and-pull model.
@@ VI-C Lookup Maps for Agent Decisions
( (a))
( (b))
Figure 15: Lookup maps for decision-making at the SA with (a) $C_{\alpha,\rm max}=0.06$ and (b) $C_{\alpha,\rm max}=0.08$.
( (a))
( (b))
Figure 16: Lookup maps for decision-making at the AA with (a) $C_{\beta,\rm max}=0.06$ and (b) $C_{\beta,\rm max}=0.08$.
As the modeled CMDPs in Section IV-B have finite states, we can depict optimal mode-based decisions derived in Algorithm 1 via a multi-dimensional lookup map for each decision-making agent. Figs. 15 and 16 illustrate the maps for the SA and the AA, respectively, under different maximum discounted costs. The number of dimensions in a map relies on the number of elements constructing every state of the relevant CMDP, with each dimension assigned to one element. With the lookup map in hand, an agent can make optimal decisions at each slot based on its current state. When comparing the same maps for different maximum discounted costs, the observation emerges that the more stringent the cost constraint is, the narrower the agent decision boundaries become. Thus, the maps could vary by changes in the parameters’ values or the goals with different target effectiveness grades. It is noteworthy that in Fig. 16, the maps with $\Theta_{n}=1$ represent the pull-based model, while the push-and-pull model converges to the push-based model with $\Theta_{n}\geq 3$ and $\Theta_{n}\geq 2$ in Figs. 16 (a) and 16 (b), respectively.
We compute an optimal threshold for each element of the state as a decision criterion, given the values of the other elements. Let us consider $\Omega_{\alpha}$ and $\Omega_{\beta}$ as the decision criteria at the SA and AA, respectively. To derive the optimal decision $\alpha_{n}^{*}$ at the $n$-th slot, there are two alternative ways to define the criterion:
@@
    $\alpha_{n}^{*}=\mathbbm{1}\big{\{}i\geq\Omega_{\alpha}(\hat{E}_{n})\,\lvert\,\hat{E}_{n}\big{\}}.$ (33)
@@
    $\alpha_{n}^{*}=\mathbbm{1}\big{\{}\hat{E}_{n}\geq\Omega_{\alpha}(v_{n})\,\lvert\,v_{n}\big{\}}.$ (34)
@@ VII Conclusion
We investigated decision-making for enhancing the effectiveness of updates communicated in the end-to-end status update system based on a push-and-pull communication model. To this end, we considered that the SA observes an information source and generates updates, and its transmission controller decides whether to send them to the AA or not. On the other hand, the AA is responsible for acting based on the updates that are successfully received and the queries initiated to accomplish a subscribed goal at the endpoint. After defining the GoE metric, we formulated the decision problem of finding optimal effect-aware policies that maximize the expected discounted sum of the update’s effectiveness subject to cost constraints. Using the dual problem, we cast it to a CMDP solved separately for each agent based on different model components and proposed an iterative algorithm to obtain the decision policies. Our results established that the push-and-pull model, on average, outperforms the push- and pull-based models in terms of energy efficiency and effectiveness, respectively. Furthermore, effect-aware policies at both agents significantly enhance the effectiveness of updates with a considerable difference in comparison to those of periodic and probabilistic effect-agnostic policies used at either or both agent(s). Finally, we proposed a threshold-based decision policy complemented by a tailored lookup map for each agent that employs effect-aware policies. Future works could explore multi-SA scenarios with varying sensing abilities and realizations’ time-dependent importance.